When buying an item at the store, you receive $9.25 back in change from the $30 you gave the cashier. If the tax rate t is 5.5%, what is the pre-tax price p? Round off to the nearest cent.
1
Expert's answer
2020-10-08T12:09:47-0400

Solution:


"Pre-tax\t \\;price \\;(p) = TP-[(\\frac{TP}{1+r} ) \\times r]"


"Where \;TP=Total \;Price \;paid"


"r= Tax\; rate"


"TP=\$30-\$9.25=\$20.75"

"r=5.5\%"


"p=20.75-[(\frac{20.75}{1+0.055} ) \times 0.055]"


"p=20.75-1.08"


"p=19.67"


"Pre-tax\; price (p)=\$19.67"